1. A closed flask contains gas at STP (273K, 76 cmHg). If the temperature is raised to 400K what will the pressure be in mmHg?
1
Expert's answer
2021-09-10T00:41:53-0400

At STP


"\frac{P_1}{T_1} = \frac{P_2}{T_2}"


P1 = 76cmHg , T1 = 273K ,P2 = ? , T2 = 400K


"\frac{76cmHg}{273K} = \frac{P_2}{400K}"


"P_2 = \frac{400K×76cmHg}{273K}"


"P_2 = 111.35cmHg"